CH/Doc/92/06

Charissa

Physics Meeting

To be held on: 4 November 1992

To be held at: Daresbury

A one day meeting is being arranged to review the Charissa physics programme. Hopefully this will give us all a chance to catch up with where the analysis of recent experiments has got to, to find out about the recent work on the cluster model, and to have a discussion about how the physics is developing and what will form the next phase of experiments that is needed.

Programme

1100  ^{24}Mg breakup off ^{16}O target  Chris Jones
1120  Higher excitations in ^{24}Mg breakup  Mike Leddy
1140  Energy dependence of ^{24}Mg breakup  Godfrey Gyapong
1200  ^{16}O+^{8}Be breakup with ^{24}Mg and ^{20}Ne  Jo Pople
1220  LUNCH
1330  Breakup in ^{24}Mg+^{24}Mg reaction  Alex Murphy
1350  Chain states in ^{16}O+^{12}C  Peter Simmons
1410  Discussion on replay and analysis needs
1430  Developments in cluster calculations  Alan Merchant
1545  Present position and future directions  Bill Rae
1600  Open discussion

B.R. Fulton W.N. Catford (Secretary) 25 August 1992